Show that y=c1epower x +c2e power 2x is the general solution of y"-3y'+2y=0 on any interval and find the particular solution for which y(0) =1




1
Expert's answer
2022-02-06T16:20:23-0500

We want to show that "\displaystyle 
y=c_1e^x+c_2e^{2x} \text{ is the general solution of }y\prime\prime-3y\prime+2y=0\\" on any interval and find the particular solution for which "\displaystyle
y(0)=1".

Now,

"y=c_1e^x+c_2e^{2x}\\
y\prime=c_1e^x+2c_2e^{2x}\\
y\prime\prime=c_1+4c_2e^{2x}"

Substituting "\displaystyle
y, y\prime, y\prime\prime" into the given DE yields;

"\displaystyle
(c_1e^x+4c_2e^{2x})-3(c_1e^x+2c_2e^{2x})+2(c_1e^x+c_2e^{2x})=0\\
\Rightarrow c_1e^x+4c_2e^{2x}-3c_1e^x-6c_2e^{2x}+2c_1e^x+2c_2e^{2x}=0\\
\Rightarrow e^x(c_1-3c_1+2c_1)+e^{2x}(4c_2-6c_2+2c_2)=0\\
\Rightarrow e^x(0)+e^{2x}(0)=0\\
\Rightarrow 0=0"

Hence, we have shown that "\displaystyle 
y=c_1e^x+c_2e^{2x}" is the general solution of the given DE on any interval.


Next, at "\displaystyle
y(0)=1", we have;

"\displaystyle
1=c_1+c_2\Rightarrow c_1=1-c_2"

Thus, the particular solution is for which "\displaystyle
y(0)=1" is; "\displaystyle 
y=c_1e^x+c_2e^{2x}=(1-c_2)e^x+c_2e^{2x}"